24.03 Paper on Conceptual Schemes

Matthew Gray

April 9, 1996

Introduction

This paper will discuss some of the formulations of conceptual schemes and try to evaluate wether these formulations are coherent. Primarily, the focus will be on Davidson’s notion and examination of conceptual schemes and some serious potential problems that arise from this formulation.

Davidson’s Conceptual Schemes

Generically, a “conceptual scheme” is a complete set of concepts, a complete view of the world. Kuhn’s paradigms could be considered one form of conceptual scheme, while others, like Davidson, take a less subtle definition. Davidson presents the notion of a conceptual scheme which is reflected by and wholly embodied by a language. While this point itself may be the subject of some contention, it is first necassary to accept Davidson’s equivocation of a “conceptual scheme” and a “language”.^{1} Given that a language is (according to Davidson) a complete conceptual scheme, and that any language intertranslatable with English shares the same conceptual scheme, Davidson generalizes the conceptual scheme of English speakers to be that of all speakers of natural language.

In fact, Davidson further expands on this notion to claim that it is impossible to have a conceptual scheme (and therefore, a language) which is not wholly translatable into English. Davidson seems to admit the possibility that there could be disjoint conceptual schemes which are subsets of the conceptual scheme associated with English, but does not seriously explore this.

An Argument Against Davidson’s View

There are two pricniples which form a basis for a possible argument against Davidson’s view that there are no conceptual schemes distinct from our own. Those principles are:

If we can clearly imagine that p, then this is good (but not conclusive) reason to think that it is possible that p.

and If we can clearly imagine having good evidence for the proposition bf p, then this is good (but not conclusive) reason to think that the proposition that p is possible.

1Though, more specifically, Davidson does not seem to want it to be a one to one mapping, but a one to many mapping, where a “conceptual scheme” is in fact an equivalence class of languages under the equivalence relation “is intertranslatable with”.

This yields an argument which shall involve the following symbols:

• E is the conceptual scheme the corresponds to our own, as embodied by English.

• S is the conceptual scheme of a hypothetical set of beings dubbed the “Saturnians”

• P is the conceptual scheme of a hypothetical set of beings dubbed the “Plutonians”

The argument presents a proposition x that E is a subset of S and that P is a subset of S and that E and P are wholly disjoint. It seems reasonable that humans could encounter the Saturnians who could converse with humans without (conceptual) difficulties after all the translation issues were resolved.

This argument is presented as a scenario in which humans encounter some Saturnians who do in fact engage humans in translatable conversation. These Saturnians would explain to us that they also knew of beings, the Plutonians, with whom they can also converse fluently, but the Plutonians can not converse with humans at all. Hence, E and P are disjoint subsets of S.

This seems to allow for two disjoint conceptual schemes (E and P) which satisfy the above principles, in that the scenario described in the previous paragraph can be “clearly imagined” and we therefore have good reason to think that it is possible.

The Nature of Disjoint Schemes

Davidson argues agains the case of complete failure of translatability by debunking some specific examples, such as, Whorf’s claim that the Hopi language contains untranslatable metaphysical concepts which Whorf proceeds to describe in English. Davidson’s case here seems to be strong, but it does not eliminate the possibility that there are in fact schemes that have a complete failure of translation into E.

The problem with such a scheme is it is inherently indescribable, in English. A scenario which would at least present a plausible nature for such an alternate scheme follows:

There is an experience e, which, if experienced enlightens one to a conceptual scheme (N) wholly disjoint from our standard scheme, E. The nature of this experience is, of course, indescribable, just as explaining the experience of “opening ones eyes” may be impossible to relate to a blind man.

Failure of Evidence

One of the fundamental problems with this seeming argument against Davidson is that it requires that one “clearly imagine that there exist concepts which are unimaginable in our conceptual scheme”. The question of whether one can clearly imagine that there are things that are unimaginable, is dubious, at best.

Davidson, it would seem, would want to discount the ability to concieve of such things, or at least place them outside the realm of the “clearly imagined”. More important, Davidson most definitely wants to claim, that even given the scenario above describing experience e, if there were beings which held only the conceptual scheme N, that any beahavior of theirs would not in fact be interpreted as reflecting an alternative conceptual scheme, but as simple incoherence of their behavior. Hence, there would not be evidence for distinct conceptual schemes, only evidence for incoherent behavior.

Davidson’s reaction to the possibility of a human individual who has experienced e and therefore understands both schemes E and N is unclear. Most likely, he would want to claim that either the individual could, in fact, describe the things in N in English, or that if they couldn’t that they were simply being incoherent.

Other Cases For Alternative Conceptual Schemes

There seems to be a great amount of concrete evidence for concepts which are fully out of the bounds of those expressible in English. For example, though it is certainly possible to describe the process of riding a bicycle in English, there certainly seems to be some concepts associated with this that are not describable, as someone is unlikely to learn to ride a bike by mere instruction in English.

Further, there seems to be a great amount expressed in body language and facial nuance that cannot be translated into word. Additionally, in a great many physical activities such as dance, or martial arts, there are concepts learned such as “frame” and “ki” which are not completely describable in English yet can be used to communicate and require experience to fully grasp.

One might counter-argue that these are in fact not “concepts”, but “skills” in the case of the bicycling and other physical arts and “feelings” in the case of facial expressions. This argument appears weak as it seems difficult to formulate a definition of “concept” which does not include any of these things but includes most things that we would qualify as “concepts”.

Another counter-argument to the above cases might be to claim that the conceptual scheme E in fact includes all forms of communication of an English speaker, not just the verbal. This is contrary to Davidson, but perhaps a reasonable admission. It is however, weakened by the fact that in order to deny the existence of alternative conceptual schemes, the definition has to be further extended to include not only all forms of communication of an English speaker, but all the forms of communication that an English speaker might learn.^{2} This, however, threatens to degenerate into a discussion of the capacity of the human brain, such as discussing whether or not it has the capacity to conceive of large cardinals.

Using again Davidson’s formulation of conceptual schemes, one could counter the counter-Davidson argument^{3} by stating that though there is a set of concepts N, it does not constitute a conceptual scheme which a being could actually hold because there are certain fundamental concepts (perhaps such as object identity and infinitive forms) contained in a minimal conceptual scheme ϵ which are contained in all other conceptual schemes. That is, there is a set ϵ (comprised of concepts ϵ_{1}, ϵ_{2}, ϵ_{3}, ...) which is a subset of all conceptual schemes X (comprised of concepts ϵ_{1}, ϵ_{2}, ϵ_{3}, ..., x_{1}, x_{2}, x_{3}, ...).

2If not, are these non-verbalized conceptss learned by some individuals, but not others, indications of alternative conceptual schemes?

Conclusion

Conceptual schemes, as Davidson initially claims, seem to be fraught with peril. It is indeed difficult to improve intelligibility of conceptual relativism through conceptual schemes, but Davidson’s argument against alternative conceptual schemes has similar problems of intelligibility.

It is the opinion of the author that both Davidson’s argument and it’s counter-argument are incoherent by fault of the equivocation of language and conceptual schemes. However, given this requirement, the counter-argument seems a strong one and Davidson’s response alleging the failure of the evidence is weak.

3Though the result is not in accordance with Davidson.